# A. I. Bolotnikov, A. A. Irmatov

### v1

@@ Introduction
Conversely, let $P=(e_{i_{1}},e_{i_{2}},...,e_{i_{2\cdot(m+n)}})$ be a path, such that $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{2\cdot(m+n)-1}W(e_{i_{2\cdot(m+n)}})=0$. The current goal is to choose values $\alpha_{1},\alpha_{2},...,\alpha_{n}$ for $x_{1},x_{2},x_{3},...,x_{n}$, such that all clauses $u_{1},u_{2},...,u_{m}$ are satisfied. For $1\leq i\leq n$, let $\alpha_{i}=True$, if $e_{i_{2\cdot i-1}}=a_{i}$, and let $\alpha_{i}=False$, if $e_{i_{2\cdot i-1}}=d_{i}$. The next step is to check if all clauses from U are satisfied. Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause. If $e_{i_{2\cdot n+2\cdot j-1}}\in\{(P_{j},R_{j,1}),(P_{j},R_{j,2}),(P_{j},R_{j,3}),(P_{j},R_{j,4})\}$, then because $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-....+(-1)^{2\cdot(m+n)-1}W(e_{i_{2\cdot(m+n)}})=0$, either $u_{j}$ contains $x_{\alpha}$ and $e_{2\cdot\alpha-1}=a_{\alpha}$ , which means that $x_{\alpha}=True$ and $u_{j}$ is satisfied, or $u_{j}$ contains $\neg x_{\alpha}$ and $e_{2\cdot\alpha-1}=d_{\alpha}$, which means that $x_{\alpha}=False$ and $u_{j}$ is satisfied again. In both cases $u_{j}$ is satisfied after values $\alpha_{1},\alpha_{2},...,\alpha_{n}$ are assigned. Cases when $e_{i_{2\cdot n+2\cdot j-1}}\in\{(P_{j},R_{j,5}),(P_{j},R_{j,6}),(P_{j},R_{j,7})\}$ can be resolved with similar reasoning, with $x_{\beta}$ or $x_{\gamma}$ being considered instead of $x_{\alpha}$. Therefore, after values $\alpha_{1},\alpha_{2},...,\alpha_{n}$ were chosen, all clauses $u_{1},u_{2},...,u_{m}$ are satified.
@@
Therefore, this process of obtaining an improper weight function problem from a 3-SAT problem is a polynomial transformation of a 3-SAT problem into an improper weight function problem, which makes an improper weight function problem an NP-complete one.
@@
For an arbitrary graph $G(V,E)$ without loops and parallel edges, a sufficiently large prime number $p$ and an arbitrary natural $k$, $\chi_{MA(G)}(p^{k})$ is equal to the number of proper weight functions $W:E\rightarrow F_{p}^{k}$.
@@
Let $A$ be an arbitrary arrangement of hyperplanes in $R^{n}$ such that the equation of each hyperplane $A$ has integer coefficients. Let $A_{q}$ be a subset of all elements $F_{q}^{n}$ satisfying the equation of at least one of the hyperplanes $A$ modulo $p$, where $q=p^{k}$ is a prime power. The finite field method allows us to find the value of the characteristic polynomial of the arrangement $A$ at the point $q$ for sufficiently large $p$: $\chi_{A}(q)=|F_{q}^{n}\setminus A_{q}|$.
Let $q=p^{k}$ be a power of a prime number, and let $d$ be an irreducible polynomial of degree $k$ over $F_{p}$. Then there is a one-to-one correspondence between $F_{p}^{k}$ and $F_{q}$ that preserves addition and multiplication by elements of $F_{p}$: each element $(a_{1},a_{2},...,a_{k})$ from $F_{p}^{k}$ is associated with the element $a_{1}+a_{2}x+a_{3}x^{2}+....+a_{k}x^{k-1}$ in $F_{p}(x)/d\cong F_{q}$. This map induces a map from the set of functions $W:E\rightarrow F_{p}^{k}$ to the set of functions $\hat{W}:E\rightarrow F_{q}$. A proper function $W$ will be mappped to a function $\hat{W}$, which has the following property: $(\hat{W}(e_{1}),\hat{W}(e_{2}),..,\hat{W}(e_{n}))$ belongs to $F_{q}^{n}\setminus MA(G)_{q}$, due to (1) and the definition of $MA(G)$.
@@
This chapter contains a description of a cryptosystem that uses an alternating weighted path problem, which is an extention of the improper weight function problem. Given a graph $G(V,E)$, a prime number $p$, a natural number $k$, a weight function $W:E\rightarrow F_{p}^{k}$ and a vector $(a_{1},...,a_{k})\in F_{p}^{k}$, the problem is to determine, if there is a sequence of edges $(e_{i_{1}},...,e_{i_{d}})$ that form a simple path or a simple even cycle in graph $G$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{k})$. Since we can use $(e_{i_{1}},...,e_{i_{d}})$ as a clue that would allow to solve this problem in polynomial time, the problem is NP. Moreover, we can reduce a general improper weight function problem to a specific alternating weighted path problem by picking $G,p,k,W$ as they were in the initial improper weight function problem and choosing $(a_{1},...,a_{k})={\bf 0}\in F_{p}^{k}$, making alternating weighted path problem an NP-hard problem. Since alternating weighted path problem is both NP and NP-hard, it’s NP-complete.
@@
Despite the general alternating weighted path problem being NP-complete, there are special cases of this problem that can be solved fairly quickly. For example, let G be a complete graph $K_{m}$, let p be a large prime, $p>2\cdot 3^{m-1}$, and let $k=m$. Let $V=\{v_{1},...,v_{m}\}$ be a set of vertices of graph G. For an edge $e=(v_{q},v_{r}),q<r$, let the weight function $W(e)=(w_{1}(e),...,w_{m}(e))$ be defined the following way: $w_{t}(e)=0$, if $t\neq q,t\neq r$, $w_{t}(e)=3^{r-q-1}$, if $t=q$, and $w_{t}(e)=1$, if $t=r$. The key idea of the solution will be for a given $x\in F_{p}$ to either find an $1\leq i\leq m$ such that $x\equiv 3^{i-1}(mod\leavevmode\nobreak\ p)$ or $x\equiv p-3^{i-1}(mod\leavevmode\nobreak\ p)$, or find $1\leq i<j\leq m$, such that $x\equiv 3^{j}-3^{i}(mod\leavevmode\nobreak\ p)$ or $x\equiv p-3^{j}+3^{i}(mod\leavevmode\nobreak\ p)$. Note that because $p>2\cdot 3^{m-1}$, and because for any $1\leq i<j\leq m$ and $r\in\{0,1,3,...,3^{i-1}\},s\in\{0,1,3,...,3^{j-1}\}$ $3^{j}-s>3^{i}-r$, it is impossible to get multiple different answers for an $x$ as a result of this search. Also note, that since |$|\{3^{i-1}|1\leq i\leq m\}\cup\{p-3^{i-1}|1\leq i\leq m\}\cup\{3^{j}-3^{i}|1\leq i<j\leq m\}\cup\{p-3^{j}+3^{i}|1\leq i<j\leq m\}|=m(m+1)$, this search can be done in polynomial time.
@@
For a vector $A=(a_{1},...,a_{m})\in F_{p}^{k}$, there is the following solution to this problem. The solution consists of m-1 steps, and the i-th step ends with either the conclusion, that the answer is “No”, or with the set of the edges $(v_{i},v_{j}),i<j$ that belong to a path or an even cycle $D=(e_{i_{1}},...,e_{i_{d}})$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{m})$. In essense, this algoritm considers vertices of graph G one by one, from $v_{1}$ to $v_{m-1}$, tries to determine for each vertex, if it has adjacent edges that are part of $D$, and then adjusts the vector of weights, so that weights on edges of verices with lower indexes no longer affect the search of edges on D, adjacent to vertices with higher indexes.
@@
Let’s describe the first step. If there is a path or an even cycle $D=(e_{i_{1}},...,e_{i_{d}})$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{m})$, then there are 3 possible options. The first option is that $a_{1}=0$. In this case $D$ doesn’t contain any edges that are adjacent to $v_{1}$. The second option is that $a_{1}=3^{b},b<m$ or $a_{1}\equiv p-3^{b}$ (mod p). In this case $(v_{1},v_{b+2})$ is in $D$ and it is the only edge in $D$ that is adjacent to $v_{1}$. The third option is that $a_{1}\equiv 3^{b}-3^{c}$ (mod p), $b<m,c<m$. In this case both edges $(v_{1},v_{b+2})$ and $(v_{1},v_{c+2})$ are in $D$ and they are the only edges in $D$ that are adjacent to $v_{1}$. If $a_{1}$ doesn’t fit into any of these three options, then the answer to the alternating weighted path problem is “No”. Otherwise, the set of edges that are adjacent to $v_{1}$ and are in the path D, is stored, and the new vector of weights $B_{1}$ is obtained from $A$ in the following way. If $a_{1}=0$ then the stored set of edges is empty and $B_{1}=A$ If $a_{1}=3^{i}$ then the edge $(v_{1},v_{i})$ is stored and $B_{1}=A-W((v_{1},v_{i}))$. If $a_{1}\equiv p-3^{i}$ (mod p), then the edge $(v_{1},v_{i})$ is stored and $B_{1}=A+W((v_{1},v_{i}))$. If $a_{1}\equiv 3^{i}-3^{j}$ (mod p), then both edges $(v_{1},v_{i})$ and $(v_{1},v_{j})$ are stored and $B_{1}=A-W((v_{1},v_{i}))+W((v_{1},v_{j}))$. And this point the first step ends.
@@
For the i-th step the process is the same, except $v_{i}$ is used instead of $v_{1}$, vector $B_{i-1}=(b_{1},...,b_{m})$, that we obtained at the end of $(i-1)$-th step, is used instead of A, $b_{i}$ is used instead of $a_{1}$, and a new vector $B_{i}$ is obtained as a result instead of $B_{1}$.
@@
Now let’s describe a cryptosystem based on the alternating weighted path problem.
@@
The important part of this cryptosystem is the superincreasing sequence in the easy version of the alternating weighted path problem. Similar approach is used in the Merkle-Hellman cryptosystem that is based on the knapsack problem [10]. In the Merkle-Hellman cryptosystem, the superincreasing sequence is disguised with a modular multiplication. This way of disguise is known to be vulnerable to attacks [11], [12]. In our cryptosystem, the superincreasing sequence is disguised with a linear operator, which is a broader method of disguise than the modular multiplication. Thus there might be a variation of this cryptosystem that is less vulnerable to attacks than knapsack-based cryptosystems.

### v3

@@ Introduction
The current goal is to choose values $\alpha_{1},...,\alpha_{n}$ for $x_{1},...,x_{n}$, such that all clauses $u_{1},...,u_{m}$ are satisfied. For $1\leq i\leq n$, let $\alpha_{i}=True$, if $e_{i_{2i-1}}=a_{i}$, and let $\alpha_{i}=False$, if $e_{i_{2i-1}}=d_{i}$. The next step is to check if all clauses from $U$ are satisfied. Let $u_{j}=(t_{\alpha},t_{\beta},t_{\gamma})$ be a clause. If $e_{i_{2n+2j-1}}\in\{(P_{j},R_{j,1}),(P_{j},R_{j,2}),(P_{j},R_{j,3}),(P_{j},R_{j,4})\}$, then because $W(e_{i_{1}})-W(e_{i_{2}})+W(e_{i_{3}})-...+(-1)^{2(m+n)-1}W(e_{i_{2(m+n)}})=0$, either $u_{j}$ contains $x_{\alpha}$ and $e_{2\alpha-1}=a_{\alpha}$ , which means that $x_{\alpha}=True$ and $u_{j}$ is satisfied, or $u_{j}$ contains $\neg x_{\alpha}$ and $e_{2\alpha-1}=d_{\alpha}$, which means that $x_{\alpha}=False$ and $u_{j}$ is satisfied again. In both cases, $u_{j}$ is satisfied after values $\alpha_{1},...,\alpha_{n}$ are assigned. Cases when $e_{i_{2n+2j-1}}\in\{(P_{j},R_{j,5}),(P_{j},R_{j,6}),(P_{j},R_{j,7})\}$ can be resolved with similar reasoning, with $x_{\beta}$ or $x_{\gamma}$ being considered instead of $x_{\alpha}$. Therefore, after values $\alpha_{1},...,\alpha_{n}$ were chosen, all clauses $u_{1},...,u_{m}$ are satisfied.
@@
Therefore, this process of obtaining an improper weight function problem from a 3-SAT problem is a polynomial transformation of a 3-SAT problem into an improper weight function problem, which makes an improper weight function problem an NP-complete one. ∎
@@
For an arbitrary graph $G(V,E)$ without loops and multiple edges, a sufficiently large prime number $p$ and an arbitrary natural $k$, $\chi_{MA(G)}(p^{k})$ is equal to the number of proper weight functions $W:E\rightarrow F_{p}^{k}$.
@@
Let $A$ be an arbitrary arrangement of hyperplanes in $\mathbb{R}^{n}$, such that the equation of each hyperplane $A$ has integer coefficients. Let $A_{q}$ be a subset of all elements $F_{q}^{n}$ satisfying the equation of at least one of the hyperplanes $A$ modulo $p$, where $q=p^{k}$ is a prime power. The finite field method allows us to find the value of the characteristic polynomial of the arrangement $A$ at the point $q$ for sufficiently large $p$: $\chi_{A}(q)=|F_{q}^{n}\setminus A_{q}|$.
Let $q=p^{k}$ be a power of a prime number, and let $d$ be an irreducible polynomial of degree $k$ over $F_{p}$. Then, there is a one-to-one correspondence between $F_{p}^{k}$ and $F_{q}$ that preserves addition and multiplication by elements of $F_{p}$: each element $(a_{1},...,a_{k})$ from $F_{p}^{k}$ is associated with the element $a_{1}+a_{2}x+a_{3}x^{2}+...+a_{k}x^{k-1}$ in $F_{p}(x)/d\cong F_{q}$. This map induces a map from the set of functions $W:E\rightarrow F_{p}^{k}$ to the set of functions $\hat{W}:E\rightarrow F_{q}$. A proper function $W$ will be mapped to a function $\hat{W}$, which has the following property: $(\hat{W}(e_{1}),...,\hat{W}(e_{n}))$ belongs to $F_{q}^{n}\setminus MA(G)_{q}$, due to (1) and the definition of $MA(G)$.
@@
This section contains a description of a cryptosystem that uses an alternating weighted path problem, which is an extention of the improper weight function problem. Given a graph $G(V,E),|E|=n$, a prime number $p$, a natural number $k$, a weight function $W:E\rightarrow F_{p}^{k}$ and a vector $(a_{1},...,a_{k})\in F_{p}^{k}$, the problem is to determine, if there is a sequence of edges $(e_{i_{1}},...,e_{i_{d}})$ that form a simple path or a simple even cycle in graph $G$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{k})$. Since we can use $(e_{i_{1}},...,e_{i_{d}})$ as a clue that would allow to solve this problem in polynomial time, the problem is NP. Moreover, we can reduce a general improper weight function problem to a specific alternating weighted path problem by picking $G,p,k,W$ as they were in the initial improper weight function problem and choosing $(a_{1},...,a_{k})={\bf 0}\in F_{p}^{k}$, making alternating weighted path problem an NP-hard problem. Since alternating weighted path problem is both NP and NP-hard, it’s NP-complete.
@@
Despite the general alternating weighted path problem being NP-complete, there are special cases of this problem that can be solved fairly quickly. For example, let $G$ be a complete graph $K_{m}$, let $p$ be a large prime, $p>2\cdot 3^{m-1}$, and let $k=m$. Let $V=\{v_{1},...,v_{m}\}$ be a set of vertices of graph $G$. For an edge $e=(v_{q},v_{r}),q<r$, let the weight function $W(e)=(w_{1}(e),...,w_{m}(e))$ be defined the following way: $w_{t}(e)=0$, if $t\neq q,t\neq r$, $w_{t}(e)=3^{r-q-1}$, if $t=q$, and $w_{t}(e)=1$, if $t=r$. The key idea of the solution will be for a given $x\in F_{p}$ to either find an $1\leq i\leq m$ such that $x\equiv 3^{i-1}$(mod $p$) or $x\equiv p-3^{i-1}$ (mod $p$), or find $1\leq i<j\leq m$, such that $x\equiv 3^{j}-3^{i}$(mod $p$) or $x\equiv p-3^{j}+3^{i}$ (mod $p$). Note that because $p>2\cdot 3^{m-1}$, and because for any $1\leq i<j\leq m$ and $r\in\{0,1,3,...,3^{i-1}\},s\in\{0,1,3,...,3^{j-1}\}$ $3^{j}-s>3^{i}-r$, it is impossible to get multiple different answers for an $x$ as a result of this search. Also note, that since |$|\{3^{i-1}|1\leq i\leq m\}\cup\{p-3^{i-1}|1\leq i\leq m\}\cup\{3^{j}-3^{i}|1\leq i<j\leq m\}\cup\{p-3^{j}+3^{i}|1\leq i<j\leq m\}|=m(m+1)$, this search can be done in polynomial time.
@@
For a vector $A=(a_{1},...,a_{m})\in F_{p}^{k}$, there is the following solution to this problem. The solution consists of $m-1$ steps, and the $i$th step ends with either the conclusion, that the answer is “No”, or with the set of the edges $(v_{i},v_{j}),i<j$ that belong to a path or an even cycle $D=(e_{i_{1}},...,e_{i_{d}})$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{m})$. In essense, this algoritm considers vertices of graph $G$ one by one, from $v_{1}$ to $v_{m-1}$, tries to determine for each vertex, if it has adjacent edges that are part of $D$, and then adjusts the vector of weights, so that weights on edges of vertices with lower indexes no longer affect the search of edges on $D$, adjacent to vertices with higher indexes.
@@
Let us describe the first step. If there is a path or an even cycle $D=(e_{i_{1}},...,e_{i_{d}})$, such that $W(e_{i_{1}})-W(e_{i_{2}})+...+(-1)^{d+1}W(e_{i_{d}})=(a_{1},...,a_{m})$, then there are 3 possible options. The first option is that $a_{1}=0$. In this case $D$ doesn’t contain any edges that are adjacent to $v_{1}$. The second option is that $a_{1}=3^{b},b<m$ or $a_{1}\equiv p-3^{b}$ (mod $p$). In this case, $(v_{1},v_{b+2})$ is in $D$ and it is the only edge in $D$ that is adjacent to $v_{1}$. The third option is that $a_{1}\equiv 3^{b}-3^{c}$ (mod $p$), $b<m,c<m$. In this case, both edges $(v_{1},v_{b+2})$ and $(v_{1},v_{c+2})$ are in $D$ and they are the only edges in $D$ that are adjacent to $v_{1}$. If $a_{1}$ doesn’t fit into any of these three options, then the answer to the alternating weighted path problem is “No”. Otherwise, the set of edges that are adjacent to $v_{1}$ and are in the path $D$, is stored, and the new vector of weights $B_{1}$ is obtained from $A$ in the following way. If $a_{1}=0$ then the stored set of edges is empty and $B_{1}=A$. If $a_{1}=3^{i}$ then the edge $(v_{1},v_{i})$ is stored and $B_{1}=A-W((v_{1},v_{i}))$. If $a_{1}\equiv p-3^{i}$ (mod $p$), then the edge $(v_{1},v_{i})$ is stored and $B_{1}=A+W((v_{1},v_{i}))$. If $a_{1}\equiv 3^{i}-3^{j}$ (mod $p$), then both edges $(v_{1},v_{i})$ and $(v_{1},v_{j})$ are stored and $B_{1}=A-W((v_{1},v_{i}))+W((v_{1},v_{j}))$. And this point the first step ends.
@@
For the $i$th step the process is the same, except $v_{i}$ is used instead of $v_{1}$, vector $B_{i-1}=(b_{1},...,b_{m})$, that we obtained at the end of $(i-1)$th step, is used instead of A, $b_{i}$ is used instead of $a_{1}$, and a new vector $B_{i}$ is obtained as a result instead of $B_{1}$.
@@
Now let us describe a cryptosystem based on the alternating weighted path problem.
@@
Such a system might require some additions, so, it can be used in transmissions of different kinds of messages. For example, a map from the set of edges to the set of letters from English alphabet can be added to the public key, so, that the system can be used to transmit English words. To further explain this idea, let us add a map $t:t(a)=$ “A”, $t(b)=$ “I”, $t(c)=$“N”, $t(d)=$ “K”, $t(e)=$“O”, $t(f)=$“M” to the example above. If the sender decides to send the message “NO”, he chooses the path $(c,e)$ that corresponds to this message according to the map $t$, computes $W(c)-W(e)=(3,0,17,-12)$ (mod 23), and sends the message $A=(3,0,17,-12)$. The receiver then computes $M(A)=(9,20,0,0)$ (mod 23). On the first step of the solution of the easy alternating path problem edge $c$ is stored and vector $(0,-3,0,-1)$ is obtained. On the second step of the solution edge $e$ is stored and vector $(0,0,0,0)$ is obtained. After that, no more edges will be added through the easy alternating path problem solution, and after applying $t$ to the result $(c,e)$, the receiver obtains the message “NO”.
One of the issues here is that the sender can use a path $(e_{i_{1}},..,e_{i_{d}})$ to send either $(t(e_{i_{1}}),...,t(e_{i_{d}}))$ or $(t(e_{i_{d}}),...,t(e_{i_{1}}))$. In case when $(e_{i_{1}},..,e_{i_{d}})$, it is easy to draw a distinction between these cases: if $A=W(e_{i_{1}})-...-W(e_{i_{d}})$, then $(t(e_{i_{1}}),...,t(e_{i_{d}}))$ was sent, and if $A=W(e_{i_{d}})-...-W(e_{i_{1}})$, then $(t(e_{i_{d}}),...,t(e_{i_{1}}))$ was sent. But if the path had odd length, the encryption process would give the same result for $(t(e_{i_{1}}),...,t(e_{i_{d}}))$ and $(t(e_{i_{d}}),...,t(e_{i_{1}}))$. Cycles of even length have a similar problem: the encryption process would give the same result for $(t(e_{i_{1}}),...,t(e_{i_{d}}))$ and $(t(e_{i_{3}}),...,t(e_{i_{d}}),t(e_{i_{1}}),t(e_{i_{2}}))$. This issue can be fixed by adding a condition to the public key, that only messages based on paths of even length can be sent. Since the path constructed in the proof of NP-completeness had an even length, the addition of an even path condition shouldn’t make the cryptosystem more vulnerable to attacks.
The important part of this cryptosystem is the superincreasing sequence in the easy version of the alternating weighted path problem. Similar approach is used in the Merkle–Hellman cryptosystem that is based on the knapsack problem [10]. In the Merkle–Hellman cryptosystem, the superincreasing sequence is disguised with a modular multiplication. This way of disguise is known to be vulnerable to attacks [11] and [12]. In our cryptosystem, the superincreasing sequence is disguised with a linear operator, which is a broader method of disguise than the modular multiplication. Thus there might be a variation of this cryptosystem that is less vulnerable to attacks than knapsack-based cryptosystems.